The decay rate of iodine-131
is 9.6% per day. What is its half-life?

Answer by Alan3354(69443)

The decay rate of iodine-131
is 9.6% per day. What is its half-life?
----------------
At the end of each day, 90.4% is left (100 - 9.6) = 0.904 of the day before.
0.5 = 0.904^t (t in days)
ln(0.5) = t*ln(0.904)
t = ln(0.5)/ln(0.904)
t =~ 6.868 days
------------------
You could use log instead of ln, the results are the same.